News Alert: GOHSEP SPECIAL WEATHER REPORT
Posted: 10/7/2013 8:21:21 AM
Louisiana Business Emergency Operations Center




6 October 2013



As of 1000
Hours



 



STATE EOC ACTIVATION LEVEL: 4 Normal Operations



WebEOC Incident Name: 031-13
Tropical Storm Karen - 3 October 2013



Note:
Current H-Hour is available in WebEOC under the H-Hour board.



**Updates
to this report will be in bold red letters**



**FINAL
REPORT**



 




  1. SITUATION: As of
         1000 hours CDT, NWS/NHC has downgraded Tropical Depression Karen to a
         remnant low pressure system and has discontinued advisories. 



 



 




  1. DECLARATIONS
        



 



      State
of Louisiana



Ascension



             
Assumption



 East
Baton Rouge



Jefferson



Lafourche



Orleans



Plaquemines



St.
Bernard



St.
Charles



St.
Helena



St.
Tammany



St.
John the Baptist



Tangipahoa



 




  1. ACTIONS
         BY PARISH
    – Most current situational reports can be found in
         "Parish SITREP to SEOC" in WebEOC. All
         parishes back to normal operations.



 



Plaquemines: Parish
declared emergency 10/3/13. EOC Level III starting 4 October 2013 through the
weekend. Request for USACE LNO and LANG LNO to report to parish EOC. EOC
   became fully activated at 0800. Shelter for 200 people will
be open tonight by 1800hrs at Belle Chasse Auditorium. Evacuations: East Bank
of Parish mandatory effective 6pm; West Bank of Parish mandatory for area south
of Ironton outside levee protection and voluntary for area south of Ironton
inside levee protection.  Empire and Ostrica Locks closing at 6pm on
10/4/13. Oakville Sluice Gate will stay open until water gets higher than Ollie
Canal Braithwaite railroad gate closed; will reopen ASAP following event.  Belle
Chasse Auditorium has received approximately 83 residents for sheltering. 
A mandatory 24-hour curfew was put into effect immediately for the East Bank of
the Parish. Lifted the mandatory evacuation order, but leaves voluntary
order in place. The parish shelter will remain open and transportation
assistance will continue to be available for residents 10/5/13 1104 hours.
Dusk-till-dawn curfew in effect on the East Bank except for residents who
evacuated and are returning home. Bell Chase shelter
closed. Returning to normal operations.



 



Terrebonne: EOC Level III
as of 3 October 2013. Closure of some Morganza System canals and floodgates.
Shelter on standby. Bayou Grand Caillou and the Houma Navigation “Bubba Dove”
Floodgates in Dulac closed as of 1800 hours today. Parish has issued a
voluntary evacuation for lower areas of the Parish as a precautionary measure.
Sand and sandbags are available to residents at predetermined sites. All
floodgates are closed at this time.



 



St. John the Baptist: The
parish has placed their shelter of last resort on standby and is currently
monitoring flooding on Hwy 51. Voluntary evacuations north of I-10 to Manchac.
Declaration received 1124. Water on road at Hwy 51 Ruddock Exit.



     



Tangipahoa: Declaration
of Emergency issued 10/3/13. At 08 00hrs on 10/4/13 EOC became fully
activated. 



     



Lafourche: Parish EOC fully
activated at 0752hrs 10/4/13. Declaration received 1100hrs 10/4/13.
Mandatory evacuation for south of the Leon Theriot

Floodgates in Golden Meadow including Port Fourchon at 2pm, today, Monday,
October 4. Raceland Recreation Center will open at 5pm 10/4/13 for Grand Isle
residents and anyone south of the floodgates in vulnerable structures. Currently,
there is 1 evacuee at the shelter.  Potentially 1 special needs individual
who will be assessed for medical conditions before potentially being transported
to the shelter. Shelter closed.  Parish is
releasing National Guard and Army Corps of Engineers. EOC is going to partial
activation and returning to normal hours 8:00am October 6, 2013 unless
situation changes. Department of Public Works Field Office are currently closed
and on standby in case needed.  Lafourche Parish Government is lifting the
mandatory evacuation from the Leon Theriot Floodgates, south.



    



 Grand
Isle:
Mandatory evacuations 10/4/13 1100hrs



    



Jefferson: EOC under
partial activation as of 1900hrs 10/4/13. Declared a state of emergency 10/4/13



    



St Bernard: EOC
partially activated. Declared a state of emergency 10/5/13. Voluntary
evacuation called for residents living outside the levee protection system.
Bayou Bienvenue & Bayou Dupre locks closed until further notice.  As
of 17:32 on 10/5/13, EOC is deactivated; Crisis Action Team activated. As of 0800 10/6/13 back to normal operations.



     



St Charles: partial
activation through end of event. Declared a state of emergency 10/4/13. Normal Operations.



    



Orleans: EOC activated at
2200 hrs on 10/4/13. Declared a state of emergency 10/4/13. EOC is deactivated.
Crisis Action Team is activated 1500 hours. As of
10/6/13 back to normal operations.



    



Assumption: Declared a
state of emergency 10/4/13 1330hrs . All activity at the site of the sinkhole
has been shut down.   



     



Ascension: Ascension has
signed a Parish Declaration of Emergency 10/4/13 1158.



     



St Tammany: Declared a
state of emergency 10/4/13. Opened Two Barns to provide sandbags to the public.
As of 2200hrs 10/5/13, parish is deactivating EOC
and returning to normal operations.



     



East Baton Rouge:
Declared a state of emergency 10/4/13 1632hrs



    



St. Helena: Declared a
state of emergency 10/4/13



 



St. Mary: released
Coast Guard LNO, Sheriff LNO, and the 2 National Guard LNO from EOC. Harvest
Moon still taking place in Franklin 11/5/13.



 



    
       All other parishes are on standby and are
monitoring the situation.



 



 



4.     ACTIONS BY AGENCY – Most current situational reports can be found in
"Agency SITREP to SEOC" in WebEOC. All
agencies back to normal operations.



 



 



GOHSEP: Governor Jindal has declared a state of emergency
10/3/13. UCG meeting scheduled 1810 hrs 10/3/13. Activation of CAT at 0730
hours on 10/3/2013. Created WebEOC incident 031-13 Tropical Storm Karen - 3
October 2013



   



DCFS: activated at Alert status. Performed statewide EP
Commo Check with 700mhz radios. Verification of contact lists throughout
agency. EOC is prepared for potential operations. DCFS staff have reported to
the shelters for SA. DCFS staff will continue to visit Shelters for SA
during the day. DCFS-EP are tracking and monitoring the storm and checking in
with effected parishes.



   



DHH: Has made contact with all federal and state ESF8
partners to inform them of situation and potential storm activity. Facilities have
checked their generators and emergency supplies. Staff on standby; has
activated its readiness contract for supplemental staff support for any
potential Search and Rescue or post-landfall operations.
  As a precautionary measure, DHH will close oyster harvesting
areas 1-23, which are located from Lake Borne in St. Bernard to Oyster Bayou in
Terrebonne parish, effective at sunset on Friday, October 4. This covers from
the eastern side of the Mississippi over to just west of the Atchafalaya
Basin.  DHH EOC reports that New Orleans EOC is
deactivating. Staff will be on standby.



  



 DOTD:
The ESF-1 EOC will be staffed 24/7 starting
10-5-13 until further notice. Dist. 02 will be prepared to support search and
rescue missions with LNOs as needed.
Plans and procedures are being
reviewed. Confirming ESF1's initiatives with school bus availability and
drivers with partners DOE and LANG. ESF3 has conducted all required
internal/external comms checks. I55 service road at Ruddock – high water but
passable. LA 1077 Lake Road South of Madisonville road is closed. EOC open
through the weekend. Toll Collections for LA 1 has been suspended to
accommodate traffic flow by noon 10/4/13.  New Orleans will be shutting
down the Ostricia Lock and evacuating personnel at 1800hrs.  DOTD District
03 (Lafayette) reported that Water is a few inches from the caps at Humble
Canal on LA 3147 which may require opening the bridge to marine traffic this
evening.



   



LDAF: Incident Command team notified. LDAF PIO issuing
media advisory for agricultural, livestock and pet owners.  Notifying NGO
partners, Fuel contractor and Dixon Correctional of standby status. 
Prepare to provide parishes with resources upon request.



 



LSP: Emergency Operations
Center: Level 2 Activation. Public Affairs: Public Information Officers are
staffing the EOC Media Room and maintaining contact with media outlets and the
public through social media. 
EOC activated with minimal staff at
0600hrs 10/4/13. LSP air support is on standby for support functions. 
Troops A, B, C, and L have been placed on standby including CIMS Team members,
BOI Rapid Response Team members and all other DPS/LSP commissioned personnel.
ESF2 has completed communications responsibility preparations.



   



DOC: Evacuated inmates from Orleans Parish Prison to EHCC.



   



Red
Cross:
Delivering mass care items
to Plaquemines Parish Shelter 10/4/13. 7 shelters on standby. Currently
contacting Parish OEPs and partners. Supporting Plaquemines shelter with water
and snacks.  Distributed snacks, water, cots and blankets to Jefferson
Parish Managed Shelter – Pard Park Playground in Marrero, Louisiana and
distributed snacks and water to Plaquemines Parish Managed Shelter – Belle
Chasse Auditorium in Belle Chasse, Louisiana. Secured and are holding snacks
and water for St. John the Baptist Parish for their Last Resort Shelter. 
Tracking all independent shelters in NSS



  



National
Guard:
Camp Beauregard and GOHSEP
on standby 24/7. LNOs deployed to: Grand Isle, Jefferson, Lafourche,
Orleans, Plaquemines, St. John, St. Tammany.



  



 LA
VOAD:
  Hosted a coordination call
10/4/13 with members and partners to report status. Additional calls will be
scheduled as needed.



  



USACE: EOC activated and working 24hr operations. USACE LNOs sent
to the following parish locations: Orleans, Lafourche, Plaquemines. Technical
assistance was provided to Plaquemines Parish to facilitate the closure of Hero
Canal Stoplogs. There are a total of 493 closures within HSDRRS. 130 have been
closed. 203 are not needed for this event. The following major structures have
been closed: Lake Bourne Surge Barrier - Barge Gate and Hero Canal Stoplogs.
USACE LGLs sent to the following Jefferson Parish at 042000OCT13. There are
a total of 493 closures within HSDRRS. 146 have been closed. 203 are not needed
for this event. West Closure Complex's - Estelle Canal and Bayou Dupree have
been closed. We continue to monitor water levels throughout the area for
potential closure of the remaining structures. St. John the Baptist Parish
CANCELLED requested supplies.
Bayou
Segnette closed at 10/5/13. LGL’s to Jefferson, St. Tammany 10/5/13.  The following major structure has been closed: Harvey Sector
Gate at 05OCT13.



  



 DEQ: Incident Command has been
established and is on stand-by as of 0700 and is monitoring the storm. Parish
liaisons have been in contact with potentially effected parishes.



 



 



 



GOHSEP Operations is monitoring the situation and will continue
to update and report if needed. Information pertaining to this storm and
all important documents from GOHSEP can be found in the file
library board within WebEOC.



 



 



 





 



 



 



The LA BEOC Portal is designed and developed by the NIMSAT Institute under the Cooperative Endeavor Agreement with the Louisiana Governor's Office of Homeland Security and Emergency Preparedness (GOHSEP)

GOHSEP LED NIMSAT

© Copyright 2010-2018, NIMSAT. All rights reserved. Terms of Service. Portal Guidelines.